Minutes of the 28th SOHO SWT Meeting

IAS, Orsay, France 21 June 1999 Contents

1 Welcoming Address 3 2 Action Items 3 3 Agree Agenda and Action Revision 3 4 Spacecraft status 3 5 Gyroless Study Progress Report 3 6 Mercury transit 3 7 Instrument status 4

7.1 Golf (A. Gabriel) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 7.2 VIRGO (C. Froehlich) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 7.3 MDI (P. Scherrer) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 7.4 SUMER (P. Lemaire) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 7.5 CDS (R. Harrison) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 7.6 EIT (J.P. Delaboudiniere) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 7.7 UVCS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 7.8 SWAN (J.L. Bertaux) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 7.9 CELIAS (P. Bochsler) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 7.10 COSTEP (H. Kunow) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 7.11 ERNE (J. Torsti) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 7.12 LASCO . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

8 Innovations 5 9 Presentation on the 1999 Leonids event 5 10 Science Priorities for the Next 4 Months 6 11 Eclipse-99 Activities 6 12 PR and outreach activities 6 13 Data availability 6 14 Status of the SOHO Archive 6 15 MEDOC Campaign 6 16 Meetings and Workshops 6 17 AOB 7 Annex 1: Agenda 8 Annex 2: List of Participants 10 Annex 3: Spacecraft Status 12 Annex 4: Gyroless Study Progress Report 24 Annex 5: Mercury Transit 27 Annex 6:Instrument Status 29

SWAN (J.P. Bertaux) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 CELIAS (P. Bochsler) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 COSTEP (H. Kunow) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 GOLF (A. Gabriel) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 CDS (R. Harrison) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 SUMER (P. Lemaire) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 VIRGO (V. Domingo) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 LASCO (R. Howard) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 ERNE (J. Torsti) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

Annex 7: Innovations 70 Annex 8: Presentation on the 1999 Leonids Event 72 Annex 9: Eclipse’99 Activities 82 Annex 10: PR Activities 90 Annex 11: Status of the SOHO Archive 104 Annex 12: MEDOC Campaign 113 Annex 13: Meetings and Workshops 128

1 Welcoming Address

Bernhard Fleck welcomed the attendees.

He introduced Stefano Coledan, a science writer who was contracted by ESA to help popularize SOHO results. PI teams are encouraged to contact the Project Scientist if they have any ideas or results that may be used for press stories.

2 Action Items

Action 28-1: On European PI’s: All the instrument teams are required to check if their flight software and ground software is Y2K compatible. The US PI’s have already done this. The European PI’s are requested to report back to L. Sanchez before mid July

3 Agree Agenda and Action Revision

No changes to the agenda (see annex 1). List of attendees - annex 2

Actions Revision

Action 27-1: Open On PI’s to provide input to what the instruments plan to do during the Mercury transit.

Action 27-2: Closed On PS to provide the orbit data for Nov. 15-17

Action 27-3: Open On EIT team to provide documentation on their roll determination method by end of March.

Action 27-4: Closed On PI’s to provide ideas for press conference at the SOHO-8 before end of April.

4 Spacecraft status

J.-P. Olive presented the spacecraft status (annex 3) and described the new gyroless spacecraft mode CRP (Coarse Roll Pointing) The spacecraft will enter this mode rather than ESR and can stay in CRP for several days. MDI is used to track roll. B. Fleck asked the EIT team to implement a system to provide real-time roll info, too.

5 Gyroless Study Progress Report

J.-P. Olive presented an overview of the gyroless mode (annex 4).

6 Mercury transit

Joe Gurman informed about Mercury’s transit on November 15-16 1999. See plot in annex 5.

7 Instrument status

The PI’s presented the status of their instruments (see annex 6).

7.1 Golf (A. Gabriel)

(see annex 6)

7.2 VIRGO (C. Froehlich)

(see annex 6) Investigating degradation/calibration issues after the SOHO interruption last year.

7.3 MDI (P. Scherrer)

No representative.

7.4 SUMER (P. Lemaire)

(see annex 6) Discussed the remaining life time of detector A and B. With continuous use they estimate the detector A and B to last for about 2 and 3 months, respectively. A fallback mode is being investigated. It will utilize the lower 30 pixels of the detector, which has not yet been exposed to light. The upper unexposed part can not be reached.

A drift in solar images was reported from the SUMER team. It was in the range ± 2 arcseconds in solarX and ± 0.5 arcsec in solarY. This is not consistent with MDI for the same period (approximately 1 arcsec in X).

7.5 CDS (R. Harrison)

(see annex 6) CDS is almost fully booked until November. GIS detector 2 is not in use at the moment. Very little scientific impact since NIS is used by most people. Have plans for the upcoming eclipse, the Mercury transit and the Whole Sun Month.

CDS has been mentioned as a source of pointing disturbances for other instruments. Instrument teams should look into the period 9-14 June 1999. In this period CDS did not do any OPS movements at all so this could be a nice check.

7.6 EIT (J.P. Delaboudiniere)

No effects on the CCD was found after 4 hours of shutterless operation.

7.7 UVCS

No representative

7.8 SWAN (J.L. Bertaux)

(see annex 6) Presented results from the press release (detection of active regions on far side of the Sun + shadow of comet Hale-Bopp). A question was raised about when one can predict active regions on the far side of the Sun on a regular bases. This is a manpower problem but CNES has indicated interest to support such a study.

7.9 CELIAS (P. Bochsler)

(see annex 6)

7.10 COSTEP (H. Kunow)

(see annex 6) Noted that there were no SOHO key parameters on the CD-rom from ISTP since October 1998. L. Sanchez will investigate this.

7.11 ERNE (J. Torsti)

(see annex 6)

7.12 LASCO

(see annex 6) Test on C1 still ongoing

8 Innovations

R. Harrison presented some thoughts about how we run SOHO today and if this should be the way we run it in the future. Should there be changes now that we are in extended operation? (see annex 7)

With the new CRP mode, there might be a possibility to conduct roll maneuvers in the future. This will be discussed in more detail at the next SWT.

The use of flags was also discussed. Should EIT produce flags? It was pointed out the e.g. bright points had been very hard to detect for issuing flags. EIT and CDS will discuss this more off-line.

SUMER has requested LASCO comet flags and is ready to observe comets close to the Sun.

More use of real-time observations was discussed. This will also be discussed in more detail at the next SWT.

9 Presentation on the 1999 Leonids event

J.-P. Olive presented information about the 1999 Leonids (annex 8).

Each instrument team was asked about their plans for this year’s Leonid storm.

GOLF - No action. Continue to observe as last year. VIRGO - No action SUMER - Leave door open CDS - Close doors EIT - Close doors LASCO - Close doors CELIAS - Still open SWAN - No action, plan to observe COSTEP - Power down ERNE - Power down

10 Science Priorities for the Next 4 Months

Not discussed.

11 Eclipse-99 Activities

P. Brekke gave an overview of the SOHO eclipse activities and the ESA eclipse web resources (annex 9)

12 PR and outreach activities

P. Brekke gave an overview of the on-going an planned PR/outreach (see annex 10).

There was an animated discussion about PR and outreach activities and the lack of coordination of some of these. ESA’s role in educational outreach was also discussed.

13 Data availability

B. Fleck mentioned concerns that some colleagues in the wider solar physics community might consider SOHO a ”closed club”, despite the very open data policy of most SOHO teams. To forestall such perceptions, he asked that - in line with the data rights policy as described in the Science Operations Plan - all data until the mission interruption one year ago should be made public. A. Gabriel expressed reluctance to do so for GOLF data. After an animated discussion it was agreed to resolve this issue off line.

14 Status of the SOHO Archive

I. Scholl gave an overview of the new archive system and gave a demonstration of the new user interface (annex 11). The PI’s were asked to review these pages and request modifications before July 16.

B.F. mentioned that the SOHO server had received more than 25 million requests.

15 MEDOC Campaign

J.C. Vial summarized the last MEDOC campaign and plans for the next MEDOC campaign (annex 12). After some discussion it was agreed to have the next MEDOC campaign during the first two weeks of October (week 40 and 41), hoping that the gyroless activities will have been successfully completed by then.

16 Meetings and Workshops

B. Fleck presented a list of upcoming meetings, which is also available at the SOHO home page (see annex 13). He mentioned in particular the IAU Symposium “Recent insights into the Physics of the Sun and Heliosphere - Highlights from SOHO and other space missions”, which will be held in Manchester on 7-10 August 2000 during the next General Assembly of the IAU.

P. Bochsler announced plans for a combined SOHO-ACE workshop. The initial plan was to hold it during the first week of year 2001. Some attendees expressed difficulties with such a date. P. Bochsler agreed to try to find another date. March 12-17 was suggested. For more details see Annex 13.

H. Kunow announced an international workshop on the Origin, Development and Effects of Coronal Mass Ejections which will take place on 27 March - 1 April 200 in Garmisch-Partenkirchen, Germany. For more details see Annex 13.

17 AOB

Next SWT meeting, SWT-29, will take place on 21-22 October at GSFC. SWT-30 will be held on 23-24 March 2000 at ESTEC.

Klaus Wilhelm suggested to give awards to team members for making SOHO such success (operations and science). This idea was well received. Will be taken up by the PS.

Annex 1: Agenda

8

Annex 2: List of Participants

10

Annex 3: Spacecraft Status

12

Annex 4: Gyroless Study Progress Report

24

Annex 5: Mercury Transit

27

Annex 6: Instrument Status

SWAN (J.P. Bertaux)

CELIAS (P. Bochsler)

36

COSTEP (H. Kunow)

38

GOLF (A. Gabriel)

41

CDS (R. Harrison)

46

SUMER (P. Lemaire)

50

VIRGO (C. Froehlich)

55

LASCO (R. Howard)

59

ERNE (J. Torsti)

61

Annex 7: Innovations

70

Annex 8: Presentation on the 1999 Leonids Event

72

Annex 9: Eclipse’99 Activities

82

Annex 10: PR Activities

90

Annex 11:Status of the SOHO Archive

104

Annex 12: MEDOC Campaigns

113

Annex 13: Meetings and Workshops

128